OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(-5, 1), C(0, 6), D(5, 1) and E (0, -4) by showing that all four sides are equal. A. BD = CE = square root of 50 B. BD = CE = 50 C. BC = CD = DE = EB = 50 D. BC = CD = DE = EB = square root of 50

OpenStudy (anonymous):

please help this is my last question and i really need help best answer gets a medal

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

D, the lengths are found from square root of the sum of the squares of the x difference and the y differences, so BC=sqrt[(5-0)^2 + (1=6)^2] = sqrt(50) etc.